# Model-free policy gradient for discrete-time mean-field control

### v1

@@ 5.2.1 Proofs on Estimation Biases
$A_{1,\varepsilon}\coloneqq\frac{B_{2}d^{3}|\mathcal{A}|\sigma_{3}}{2}+\sum_{k=2}^{t-1}\binom{t-1}{k-1}\left(\frac{\varepsilon^{2k-3}}{2^{k-1}}B_{1}B_{2}^{k-1}d^{2(k-1)}d^{k}|\mathcal{A}|^{k}\sigma_{2k}+\frac{\varepsilon^{2k-2}}{2^{k}}B_{2}^{k}d^{3k}|\mathcal{A}|^{k}\sigma_{2k+1}\right).$ (5.10)
@@
$\mathbb{E}\left[\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\zeta_{t}^{\theta,\varepsilon}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta},s,X_{s}^{\theta},l_{s}^{\theta})\right]=\frac{\varepsilon^{2}}{2}\mathbb{E}\left[\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta},s,X_{s}^{\theta},l_{s}^{\theta})\sum_{\tau=0}^{t-1}U_{\tau}^{\theta}\right]$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\sum_{k=2}^{t-1}\sum_{0\leq s_{1}<\ldots<s_{k}<t}\mathbb{E}\left[\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta},s,X_{s}^{\theta},l_{s}^{\theta})\prod_{r=1}^{k}V_{s_{r}}^{\theta}\right].$
@@
$A_{2,\varepsilon}\coloneqq\sum_{k=1}^{t-1}\binom{t}{k}\frac{\varepsilon^{2k-2}}{2^{k}}CB_{2}^{k}d^{3k}|\mathcal{A}|^{k}\sigma_{2k}.$ (5.11)
@@
$=\mathbb{E}\left[\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\sum_{s=0}^{t-1}\varepsilon^{-1}\Lambda_{s}\widehat{\nabla_{\theta}l_{s}^{\theta}}R_{t}^{\theta,\varepsilon}\right]$
$=\mathbb{E}\left[\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\sum_{s=0}^{t-1}\varepsilon^{-1}\Lambda_{s}\widehat{\nabla_{\theta}l_{s}^{\theta}}\left(1+\sum_{s^{\prime}=0}^{t-1}\varepsilon\nabla_{l}\log\psi(\theta,X^{\theta}_{s^{\prime}+1},s^{\prime},X_{s^{\prime}}^{\theta},\alpha_{s^{\prime}}^{\theta},l_{s^{\prime}}^{\theta})\cdot\Lambda_{s^{\prime}}+\zeta_{t}^{\theta,\varepsilon}\right)\right].$
@@
$\leq\sum_{s=0}^{t-1}\left\|\mathbb{E}\left[\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\nabla_{l}\log\psi(\theta,X_{s+1}^{\theta},s,X_{s}^{\theta},\alpha_{s}^{\theta},l_{s}^{\theta})\right]\right\|_{\infty}E_{s}+\varepsilon\sum_{s=0}^{t-1}A_{1,\varepsilon}\left(E_{s}+\left\|\nabla_{\theta}l_{s}^{\theta}\right\|\right).$
@@
$\leq\sum_{i=1}^{d}\frac{1}{\mathbb{P}(X_{t}^{\theta}=x^{(i)})}\Bigg(tL^{\prime}_{p}\frac{\varepsilon}{2}+tA_{2,\varepsilon}\varepsilon^{2}+$
$\quad\quad\quad\sum_{s=0}^{t-1}B_{1}|\mathcal{A}|d^{\mathbf{1}_{s<t-1}}E_{s}(\theta,\varepsilon)+\varepsilon\sum_{s=0}^{t-1}A_{1,\varepsilon}\left(E_{s}(\theta,\varepsilon)+\left\|\nabla_{\theta}l_{s}^{\theta}\right\|\right)\Bigg).$
@@
$\tilde{c}_{1}$ $=\mathbf{s}^{\theta}\left(\varepsilon A_{1,\varepsilon}+B_{1}|\mathcal{A}|\right),\quad\tilde{c}_{2}=\mathbf{s}^{\theta}B_{1}|\mathcal{A}|(d-1),$
$\tilde{c}_{3}$ $=\mathbf{s}^{\theta}\left(L^{\prime}_{p}\frac{\varepsilon}{2}+A_{2,\varepsilon}\varepsilon^{2}\right),\quad\tilde{c}_{4}=\mathbf{s}^{\theta}\left(\varepsilon A_{1,\varepsilon}\frac{3C}{2(\beta-1)}\right),\quad\tilde{c}_{5}=\beta,\quad,\tilde{c}_{6}=0,$
where $\mathbf{s}^{\theta}\coloneqq\sup_{t\geq 0}\sum_{i=1}^{d}\mathbb{P}(X_{t}^{\theta}=x^{(i)})^{-1}<\infty.$
Now, with these notations let $\beta_{\varepsilon}\coloneqq\frac{1}{2}(1+\tilde{c}_{1}+\sqrt{(1+\tilde{c}_{1})^{2}+4\tilde{c}_{2}})$. Since $\varepsilon A_{1,\varepsilon}>0$, we have $\beta_{\varepsilon}>\beta=\tilde{c}_{5}$, therefore
@@
$E_{t}(\theta,\varepsilon)\leq\mathbf{s}^{\theta}\left[\frac{3L_{p}^{\prime}}{4(\beta-1)}\varepsilon+\frac{3A_{2,\varepsilon}}{2(\beta_{\varepsilon}-1)}\varepsilon^{2}\right]\left(\beta_{\varepsilon}^{t}-1\right)+\mathbf{s}^{\theta}\frac{9CA_{1,\varepsilon}\beta}{4(\beta-1)(\beta_{\varepsilon}-\beta)}\varepsilon\left(\beta_{\varepsilon}^{t}-\beta^{t}\right).$
@@
$K_{1,\varepsilon}$ $=\mathbf{s}^{\theta}\frac{9CA_{1,\varepsilon}\beta}{4(\beta-1)(\beta_{\varepsilon}-\beta)},\quad K_{2,\varepsilon}=\mathbf{s}^{\theta}\left[\frac{3L_{p}^{\prime}}{4(\beta-1)}+\frac{3A_{2,\varepsilon}}{2(\beta_{\varepsilon}-1)}\varepsilon\right],$ (5.13)
$\beta_{\varepsilon}$ $=\frac{1}{2}\left(1+\mathbf{s}^{\theta}\left(\varepsilon A_{1,\varepsilon}+B_{1}|\mathcal{A}|\right)+\sqrt{\left[1+\mathbf{s}^{\theta}\left(\varepsilon A_{1,\varepsilon}+B_{1}|\mathcal{A}|\right)\right]^{2}+4\mathbf{s}^{\theta}B_{1}|\mathcal{A}|(d-1)}\right).$
@@
$E_{t}(\theta,\varepsilon)\leq\mathbf{s}^{\theta}\frac{L^{\prime}_{p}}{2}t\varepsilon,$
@@
$\Delta_{\mathrm{RF}}(\theta)\coloneqq\left\|\mathbb{E}\left[\sum_{t=0}^{T-1}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{t}^{\theta,\varepsilon},t,Y_{t}^{\theta},l_{t}^{\theta}+\varepsilon\Lambda_{t})G_{t}^{\theta,\varepsilon}\right]-\mathrm{RF}(\theta)\right\|_{\infty},$
$\Delta_{\mathrm{M}}(\theta)\coloneqq\left\|\mathbb{E}\left[\sum_{t=0}^{T}\varepsilon^{-1}\Lambda_{t}\widehat{\nabla_{\theta}l^{\theta}_{t}}G_{t}^{\theta,\varepsilon}\right]-(\mathrm{MD}(\theta)+\mathrm{MFD}(\theta))\right\|_{\infty}.$
We first examine $\Delta_{\mathrm{RF}}(\theta)$. We proceed in a similar way to the study of $b_{t,1}^{(i)}$ in the proof of Theorem 3.2. We define $H_{t}^{\theta,\varepsilon}\coloneqq\sum_{s=t}^{T-1}\tilde{r}(Y_{s}^{\theta,\varepsilon},\alpha_{s}^{\theta,\varepsilon},l_{s}^{\theta})+\tilde{g}(Y_{T}^{\theta,\varepsilon},l_{T}^{\theta})$. Then
$\Delta_{\mathrm{RF}}(\theta)$ $\leq\Delta_{\mathrm{RF},1}(\theta)+\Delta_{\mathrm{RF},2}(\theta),$
$\Delta_{\mathrm{RF},1}(\theta)$ $\coloneqq\left\|\mathbb{E}\left[\sum_{t=0}^{T-1}\left(\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{t}^{\theta,\varepsilon},t,Y_{t}^{\theta},l_{t}^{\theta}+\varepsilon\Lambda_{t})G_{t}^{\theta,\varepsilon}-\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{t}^{\theta,\varepsilon},t,Y_{t}^{\theta,\varepsilon},l_{t}^{\theta})H_{t}^{\theta,\varepsilon}\right)\right]\right\|_{\infty},$
$\Delta_{\mathrm{RF},2}(\theta)$ $\coloneqq\left\|\mathbb{E}\left[\sum_{t=0}^{T-1}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{t}^{\theta,\varepsilon},t,Y_{t}^{\theta,\varepsilon},l_{t}^{\theta})H_{t}^{\theta,\varepsilon}\right]-\mathrm{RF}(\theta)\right\|_{\infty}.$
@@
$\Delta_{\mathrm{RF},1}(\theta)$ $\leq\sum_{t=0}^{T-1}\left(L_{p}^{\prime}(T-t+1)M_{0}+\left((T-t)L_{r}+L_{g}\right)C\right)\mathbb{E}\left[d_{\mathrm{TV}}(\mu_{t}^{\theta},\mu_{t}^{\theta,\varepsilon})\right]$
@@
where $A_{3}=\left(M_{0}L^{\prime}_{p}\frac{T^{2}+3T}{2}+CL_{r}\frac{T^{2}+T}{2}+CL_{g}T\right)/2$. For $\Delta_{\mathrm{RF},2}(\theta)$, we use yet again the likelihood ratio trick. Recall $R_{T}^{\theta,\varepsilon}$ defined in (5.7) and its expansion in (5.8). We have
$\Delta_{\mathrm{RF},2}(\theta)$ $=\left\|\mathbb{E}\left[(1-R_{T}^{\theta,\varepsilon})\sum_{t=0}^{T-1}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{t}^{\theta},t,X_{t}^{\theta},l_{t}^{\theta})G_{t}^{\theta}\right]\right\|_{\infty}$
@@
$\Delta_{\mathrm{RF},2}(\theta)=\left\|\mathbb{E}\left[\zeta_{T}^{\theta,\varepsilon}\sum_{t=0}^{T-1}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{t}^{\theta},t,X_{t}^{\theta},l_{t}^{\theta})G_{t}^{\theta}\right]\right\|_{\infty}\leq A_{4,\varepsilon}\varepsilon^{2}.$
@@
$\Delta_{\mathrm{MD}}(\theta)$ $\coloneqq\left\|\mathbb{E}\left[\sum_{t=0}^{T}\varepsilon^{-1}\Lambda_{t}\widehat{\nabla_{\theta}l^{\theta}_{t}}(G_{t}^{\theta,\varepsilon}-H_{t}^{\theta,\varepsilon})\right]-\mathrm{MD}(\theta)\right\|_{\infty},$
$\Delta_{\mathrm{MFD}}(\theta)$ $\coloneqq\left\|\mathbb{E}\left[\sum_{t=0}^{T}\varepsilon^{-1}\Lambda_{t}\widehat{\nabla_{\theta}l^{\theta}_{t}}H_{t}^{\theta,\varepsilon}\right]-\mathrm{MFD}(\theta)\right\|_{\infty}.$
@@
$=\mathbb{E}\left[\left\{\sum_{t=0}^{T}\Lambda_{t}\widehat{\nabla_{\theta}l^{\theta}_{t}}\left(\sum_{s=t}^{T-1}\nabla_{l}\tilde{r}(X_{s}^{\theta},\alpha_{s}^{\theta},l_{s}^{\theta})\cdot\Lambda_{s}+\nabla_{l}\tilde{g}(X_{T}^{\theta},l_{T}^{\theta})\cdot\Lambda_{T}\right)\right\}R_{T}^{\theta,\varepsilon}\right]$
$\quad+\frac{\varepsilon}{2}\mathbb{E}\left[\left\{\sum_{t=0}^{T}\Lambda_{t}\widehat{\nabla_{\theta}l^{\theta}_{t}}\sum_{s=t}^{T}\sum_{j,k=1}^{d}\chi^{\theta}_{s,j,k}\Lambda_{s}^{(j)}\Lambda_{s}^{(k)}\right\}R_{T}^{\theta,\varepsilon}\right].$
@@
$=\mathbb{E}\left[\sum_{t=0}^{T}\Lambda_{t}\widehat{\nabla_{\theta}l^{\theta}_{t}}\left(\sum_{s=t}^{T-1}\nabla_{l}\tilde{r}(X_{s}^{\theta},\alpha_{s}^{\theta},l_{s}^{\theta})\cdot\Lambda_{s}+\nabla_{l}\tilde{g}(X_{T}^{\theta},l_{T}^{\theta})\cdot\Lambda_{T}\right)\right]+\mathcal{O}(\varepsilon).$
@@
$=\sum_{t=0}^{T-1}\mathbb{E}\left[\nabla_{l}\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta})\right]\mathbb{E}\left[\widehat{\nabla_{\theta}l^{\theta}_{t}}\right]+\mathbb{E}\left[\nabla_{l}\tilde{g}(X_{T}^{\theta},l_{T}^{\theta})\right]\mathbb{E}\left[\widehat{\nabla_{\theta}l^{\theta}_{T}}\right$].
@@
$=\mathbb{E}\left[\sum_{t=0}^{T}\varepsilon^{-1}\Lambda_{t}\widehat{\nabla_{\theta}l_{t}^{\theta}}G_{t}^{\theta}\right]+\mathbb{E}\left[\sum_{t=0}^{T}\Lambda_{t}\widehat{\nabla_{\theta}l_{t}^{\theta}}G_{t}^{\theta}\sum_{s=0}^{T-1}\nabla_{l}\log\psi(\theta,X_{s+1}^{\theta},s,X_{s}^{\theta},\alpha_{s}^{\theta},l_{s}^{\theta})\cdot\Lambda_{s}\right]$
$\quad+\mathbb{E}\left[\sum_{t=0}^{T}\varepsilon^{-1}\Lambda_{t}\widehat{\nabla_{\theta}l_{t}^{\theta}}G_{t}^{\theta}\zeta_{T}^{\theta,\varepsilon}\right]$
$=\sum_{t=0}^{T-1}\mathbb{E}\left[\nabla_{l}\log\psi(\theta,X_{t+1}^{\theta},t,X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta})G_{t}^{\theta}\right]\mathbb{E}\left[\widehat{\nabla_{\theta}l_{t}^{\theta}}\right]+\mathcal{O}(\varepsilon).$

### v2

@@ 5.2.1 Proofs on Estimation Biases
$A_{1,\varepsilon}\coloneq\frac{B_{2}d^{3}|\mathcal{A}|\sigma_{3}}{2}+\sum_{k=2}^{t-1}\binom{t-1}{k-1}\left(\frac{\varepsilon^{2k-3}}{2^{k-1}}B_{1}B_{2}^{k-1}d^{2(k-1)}d^{k}|\mathcal{A}|^{k}\sigma_{2k}+\frac{\varepsilon^{2k-2}}{2^{k}}B_{2}^{k}d^{3k}|\mathcal{A}|^{k}\sigma_{2k+1}\right).$ (5.10)
@@
$\mathbb{E}\left[\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\zeta_{t}^{\theta,\varepsilon}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta},s,X_{s}^{\theta},l_{s}^{\theta})\right]$
$\qquad=\frac{\varepsilon^{2}}{2}\mathbb{E}\left[\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta},s,X_{s}^{\theta},l_{s}^{\theta})\sum_{\tau=0}^{t-1}U_{\tau}^{\theta}\right]$
$\qquad\quad+\sum_{k=2}^{t-1}\sum_{0\leq s_{1}<\ldots<s_{k}<t}\mathbb{E}\left[\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{s}^{\theta},s,X_{s}^{\theta},l_{s}^{\theta})\prod_{r=1}^{k}V_{s_{r}}^{\theta}\right].$
@@
$A_{2,\varepsilon}\coloneq\sum_{k=1}^{t-1}\binom{t}{k}\frac{\varepsilon^{2k-2}}{2^{k}}CB_{2}^{k}d^{3k}|\mathcal{A}|^{k}\sigma_{2k},$ (5.11)
concluding the proof of Lemma 5.3. ∎
@@
$\qquad=\mathbb{E}\left[\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\sum_{s=0}^{t-1}\varepsilon^{-1}\Lambda_{s}\widehat{\nabla_{\theta}l_{s}^{\theta}}R_{t}^{\theta,\varepsilon}\right]$
$\qquad=\mathbb{E}\left[\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\sum_{s=0}^{t-1}\varepsilon^{-1}\Lambda_{s}\widehat{\nabla_{\theta}l_{s}^{\theta}}\left(1+\sum_{s^{\prime}=0}^{t-1}\varepsilon\nabla_{l}\log\psi(\theta,X^{\theta}_{s^{\prime}+1},s^{\prime},X_{s^{\prime}}^{\theta},\alpha_{s^{\prime}}^{\theta},l_{s^{\prime}}^{\theta})\cdot\Lambda_{s^{\prime}}+\zeta_{t}^{\theta,\varepsilon}\right)\right].$
@@
$\leq\sum_{s=0}^{t-1}\left\|\mathbb{E}\left[\mathbf{1}_{X_{t}^{\theta}=x^{(i)}}\nabla_{l}\log\psi(\theta,X_{s+1}^{\theta},s,X_{s}^{\theta},\alpha_{s}^{\theta},l_{s}^{\theta})\right]\right\|_{\infty}E_{s}(\theta,\varepsilon)$
$\quad+\varepsilon\sum_{s=0}^{t-1}A_{1,\varepsilon}\left(E_{s}(\theta,\varepsilon)+\left\|\nabla_{\theta}l_{s}^{\theta}\right\|\right).$
@@
$\leq\sum_{i=1}^{d}\frac{1}{\mathbb{P}(X_{t}^{\theta}=x^{(i)})}\Bigg(tL^{\prime}_{p}\frac{\varepsilon}{2}+tA_{2,\varepsilon}\varepsilon^{2}$
$\quad+\sum_{s=0}^{t-1}B_{1}|\mathcal{A}|d^{\mathbf{1}_{s<t-1}}E_{s}(\theta,\varepsilon)+\varepsilon\sum_{s=0}^{t-1}A_{1,\varepsilon}\left(E_{s}(\theta,\varepsilon)+\left\|\nabla_{\theta}l_{s}^{\theta}\right\|\right)\Bigg).$
@@
$\tilde{c}_{1}$ $=\mathfrak{s}^{\theta}\left(A_{1,\varepsilon}\varepsilon+B_{1}|\mathcal{A}|\right),\quad\tilde{c}_{2}=\mathfrak{s}^{\theta}B_{1}|\mathcal{A}|(d-1),$
$\tilde{c}_{3}$ $=\mathfrak{s}^{\theta}\left(L^{\prime}_{p}\frac{\varepsilon}{2}+A_{2,\varepsilon}\varepsilon^{2}\right),\quad\tilde{c}_{4}=\mathfrak{s}^{\theta}\left(A_{1,\varepsilon}\varepsilon\frac{3C}{2(\beta-1)}\right),\quad\tilde{c}_{5}=\beta,\quad,\tilde{c}_{6}=0,$
where $\mathfrak{s}^{\theta}\coloneq\sup_{t\geq 0}\sum_{i=1}^{d}\mathbb{P}(X_{t}^{\theta}=x^{(i)})^{-1}<\infty.$
Now, with these notations let $\beta_{\varepsilon}\coloneq\frac{1}{2}(1+\tilde{c}_{1}+\sqrt{(1+\tilde{c}_{1})^{2}+4\tilde{c}_{2}})$. Since $A_{1,\varepsilon}\varepsilon>0$, we have $\beta_{\varepsilon}>\beta=\tilde{c}_{5}$, therefore
@@
$E_{t}(\theta,\varepsilon)\leq\mathfrak{s}^{\theta}\left[\frac{3L_{p}^{\prime}}{4(\beta-1)}\varepsilon+\frac{3A_{2,\varepsilon}}{2(\beta_{\varepsilon}-1)}\varepsilon^{2}\right]\left(\beta_{\varepsilon}^{t}-1\right)+\mathfrak{s}^{\theta}\frac{9CA_{1,\varepsilon}\beta}{4(\beta-1)(\beta_{\varepsilon}-\beta)}\varepsilon\left(\beta_{\varepsilon}^{t}-\beta^{t}\right).$
@@
$K_{1,\varepsilon}$ $=\mathfrak{s}^{\theta}\frac{9CA_{1,\varepsilon}\beta}{4(\beta-1)(\beta_{\varepsilon}-\beta)},\quad K_{2,\varepsilon}=\mathfrak{s}^{\theta}\left[\frac{3L_{p}^{\prime}}{4(\beta-1)}+\frac{3A_{2,\varepsilon}}{2(\beta_{\varepsilon}-1)}\varepsilon\right],$ (5.13)
$\beta_{\varepsilon}$ $=\frac{1}{2}\left(1+\mathfrak{s}^{\theta}\left(A_{1,\varepsilon}\varepsilon+B_{1}|\mathcal{A}|\right)+\sqrt{\left[1+\mathfrak{s}^{\theta}\left(A_{1,\varepsilon}\varepsilon+B_{1}|\mathcal{A}|\right)\right]^{2}+4\mathfrak{s}^{\theta}B_{1}|\mathcal{A}|(d-1)}\right).$
@@
$E_{t}(\theta,\varepsilon)\leq\mathfrak{s}^{\theta}\frac{L^{\prime}_{p}}{2}t\varepsilon,$
@@
$\Delta_{\mathrm{RF}}(\theta)\coloneq\left\|\mathbb{E}\left[\sum_{t=0}^{T-1}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{t}^{\theta,\varepsilon},t,Y_{t}^{\theta},l_{t}^{\theta}+\varepsilon\Lambda_{t})G_{t}^{\theta,\varepsilon}\right]-\mathrm{RF}(\theta)\right\|_{\infty},$
$\Delta_{\mathrm{M}}(\theta)\coloneq\left\|\mathbb{E}\left[\sum_{t=0}^{T}\varepsilon^{-1}\Lambda_{t}\widehat{\nabla_{\theta}l^{\theta}_{t}}G_{t}^{\theta,\varepsilon}\right]-(\mathrm{MD}(\theta)+\mathrm{MFD}(\theta))\right\|_{\infty}.$
We first examine $\Delta_{\mathrm{RF}}(\theta)$. We proceed in a similar way to the study of $b_{t,1}^{(i)}$ in the proof of Theorem 3.2. We define $H_{t}^{\theta,\varepsilon}\coloneq\sum_{s=t}^{T-1}\tilde{r}(Y_{s}^{\theta,\varepsilon},\alpha_{s}^{\theta,\varepsilon},l_{s}^{\theta})+\tilde{g}(Y_{T}^{\theta,\varepsilon},l_{T}^{\theta})$. Then
$\Delta_{\mathrm{RF}}(\theta)$ $\leq\Delta_{\mathrm{RF}}^{(1)}(\theta)+\Delta_{\mathrm{RF}}^{(2)}(\theta),$
$\Delta_{\mathrm{RF}}^{(1)}(\theta)$ $\coloneq\left\|\mathbb{E}\left[\sum_{t=0}^{T-1}\left(\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{t}^{\theta,\varepsilon},t,Y_{t}^{\theta},l_{t}^{\theta}+\varepsilon\Lambda_{t})G_{t}^{\theta,\varepsilon}-\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{t}^{\theta,\varepsilon},t,Y_{t}^{\theta,\varepsilon},l_{t}^{\theta})H_{t}^{\theta,\varepsilon}\right)\right]\right\|_{\infty},$
$\Delta_{\mathrm{RF}}^{(2)}(\theta)$ $\coloneq\left\|\mathbb{E}\left[\sum_{t=0}^{T-1}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{t}^{\theta,\varepsilon},t,Y_{t}^{\theta,\varepsilon},l_{t}^{\theta})H_{t}^{\theta,\varepsilon}\right]-\mathrm{RF}(\theta)\right\|_{\infty}.$
@@
$\Delta_{\mathrm{RF}}^{(1)}(\theta)$ $\leq\sum_{t=0}^{T-1}\left(L_{p}^{\prime}(T-t+1)M_{0}+\left((T-t)L_{r}+L_{g}\right)C\right)\mathbb{E}\left[d_{\mathrm{TV}}(\mu_{t}^{\theta},\mu_{t}^{\theta,\varepsilon})\right]$
@@
where $A_{3}=\left(M_{0}L^{\prime}_{p}\frac{T^{2}+3T}{2}+CL_{r}\frac{T^{2}+T}{2}+CL_{g}T\right)/2$. For $\Delta_{\mathrm{RF}}^{(2)}(\theta)$, we use yet again the likelihood ratio trick. Recall $R_{T}^{\theta,\varepsilon}$ defined in (5.7) and its expansion in (5.8). We have
$\Delta_{\mathrm{RF}}^{(2)}(\theta)$ $=\left\|\mathbb{E}\left[(1-R_{T}^{\theta,\varepsilon})\sum_{t=0}^{T-1}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{t}^{\theta},t,X_{t}^{\theta},l_{t}^{\theta})G_{t}^{\theta}\right]\right\|_{\infty}$
@@
$\Delta_{\mathrm{RF}}^{(2)}(\theta)=\left\|\mathbb{E}\left[\zeta_{T}^{\theta,\varepsilon}\sum_{t=0}^{T-1}\nabla_{\theta}\log\tilde{p}(\theta,\alpha_{t}^{\theta},t,X_{t}^{\theta},l_{t}^{\theta})G_{t}^{\theta}\right]\right\|_{\infty}\leq A_{4,\varepsilon}\varepsilon^{2}.$
@@
$\Delta_{\mathrm{MD}}(\theta)$ $\coloneq\left\|\mathbb{E}\left[\sum_{t=0}^{T}\varepsilon^{-1}\Lambda_{t}\widehat{\nabla_{\theta}l^{\theta}_{t}}(G_{t}^{\theta,\varepsilon}-H_{t}^{\theta,\varepsilon})\right]-\mathrm{MD}(\theta)\right\|_{\infty},$
$\Delta_{\mathrm{MFD}}(\theta)$ $\coloneq\left\|\mathbb{E}\left[\sum_{t=0}^{T}\varepsilon^{-1}\Lambda_{t}\widehat{\nabla_{\theta}l^{\theta}_{t}}H_{t}^{\theta,\varepsilon}\right]-\mathrm{MFD}(\theta)\right\|_{\infty}.$
@@
$\qquad=\mathbb{E}\left[\left\{\sum_{t=0}^{T}\Lambda_{t}\widehat{\nabla_{\theta}l^{\theta}_{t}}\left(\sum_{s=t}^{T-1}\nabla_{l}\tilde{r}(X_{s}^{\theta},\alpha_{s}^{\theta},l_{s}^{\theta})\cdot\Lambda_{s}+\nabla_{l}\tilde{g}(X_{T}^{\theta},l_{T}^{\theta})\cdot\Lambda_{T}\right)\right\}R_{T}^{\theta,\varepsilon}\right]$
$\qquad\quad+\frac{\varepsilon}{2}\mathbb{E}\left[\left\{\sum_{t=0}^{T}\Lambda_{t}\widehat{\nabla_{\theta}l^{\theta}_{t}}\sum_{s=t}^{T}\sum_{j,k=1}^{d}\chi^{\theta}_{s,j,k}\Lambda_{s}^{(j)}\Lambda_{s}^{(k)}\right\}R_{T}^{\theta,\varepsilon}\right].$
@@
$\qquad=\mathbb{E}\left[\sum_{t=0}^{T}\Lambda_{t}\widehat{\nabla_{\theta}l^{\theta}_{t}}\left(\sum_{s=t}^{T-1}\nabla_{l}\tilde{r}(X_{s}^{\theta},\alpha_{s}^{\theta},l_{s}^{\theta})\cdot\Lambda_{s}+\nabla_{l}\tilde{g}(X_{T}^{\theta},l_{T}^{\theta})\cdot\Lambda_{T}\right)\right]+\mathcal{O}(\varepsilon).$
@@
$\qquad=\sum_{t=0}^{T-1}\mathbb{E}\left[\nabla_{l}\tilde{r}(X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta})\right]\mathbb{E}\left[\widehat{\nabla_{\theta}l^{\theta}_{t}}\right]+\mathbb{E}\left[\nabla_{l}\tilde{g}(X_{T}^{\theta},l_{T}^{\theta})\right]\mathbb{E}\left[\widehat{\nabla_{\theta}l^{\theta}_{T}}\right$].
@@
$\qquad=\mathbb{E}\left[\sum_{t=0}^{T}\varepsilon^{-1}\Lambda_{t}\widehat{\nabla_{\theta}l_{t}^{\theta}}G_{t}^{\theta}\right]+\mathbb{E}\left[\sum_{t=0}^{T}\Lambda_{t}\widehat{\nabla_{\theta}l_{t}^{\theta}}G_{t}^{\theta}\sum_{s=0}^{T-1}\nabla_{l}\log\psi(\theta,X_{s+1}^{\theta},s,X_{s}^{\theta},\alpha_{s}^{\theta},l_{s}^{\theta})\cdot\Lambda_{s}\right]$
$\qquad\quad+\mathbb{E}\left[\sum_{t=0}^{T}\varepsilon^{-1}\Lambda_{t}\widehat{\nabla_{\theta}l_{t}^{\theta}}G_{t}^{\theta}\zeta_{T}^{\theta,\varepsilon}\right]$
$\qquad=\sum_{t=0}^{T-1}\mathbb{E}\left[\nabla_{l}\log\psi(\theta,X_{t+1}^{\theta},t,X_{t}^{\theta},\alpha_{t}^{\theta},l_{t}^{\theta})G_{t}^{\theta}\right]\mathbb{E}\left[\widehat{\nabla_{\theta}l_{t}^{\theta}}\right]+\mathcal{O}(\varepsilon).$